OpenStudy (anonymous):

Please help! Rewrite 4^-2=1/16 as an equivalent logarithmic equation.

jimthompson5910 (jim_thompson5910):

\[\Large 4^{-2} = \frac{1}{16}\] converts to \[\Large \log_{4}\left( \frac{1}{16}\right) = -2\]

OpenStudy (anonymous):

Thanks man!

jimthompson5910 (jim_thompson5910):

sure thing